OpenStudy (anonymous):

Complete the standard form of the equation -4/3y=3-x What is the x-intercept of the line with the equation 8x-1/3y=16

OpenStudy (anonymous):

is it: \[-\frac{ 4 }{ 3y }\] or \[-\frac{ 4 }{ 3 }y\]

OpenStudy (anonymous):

-4/3 y ( the bottom one,) y is next to the "fraction"

OpenStudy (anonymous):

Do you know what's the standard form?

OpenStudy (anonymous):

no, thats what my problem is is i keep doing them wrong when trying to convert them to standard

OpenStudy (anonymous):

Standard form: Ax+By=C Slope-Intercept form: y=mx+b

OpenStudy (anonymous):

Multiply the equation by 3 to get rid of the fraction.

OpenStudy (anonymous):

-4y=3(3-x)

OpenStudy (anonymous):

oh!that sorta make sense

OpenStudy (anonymous):

Continue from here, order them around and tell me what you get.

OpenStudy (anonymous):

ok, itll take me a while. Thank you!

OpenStudy (anonymous):

Hit me up when and if you need help with the 2nd part.

OpenStudy (anonymous):

well -4*3=-12 and 3*3= 9. I know they want the whole thing to look like (_x+_y=9) so is that how i am supposed to do that?

OpenStudy (anonymous):

@Ankh

OpenStudy (anonymous):

If you multiply a fraction by its denominator, it just removes the fraction. \[-\frac{ 4y }{ 3 } \times 3 = -4y\]

OpenStudy (anonymous):

I think less advanced math call it LCM.

OpenStudy (anonymous):